OpenStudy (lovelycharm):

What is the solution to the following system of equations? y = x2 + 10x + 11 y = x2 + x − 7

OpenStudy (john_es):

I would use the fact that both equations have their left sides equal, \[x^2 + 10x + 11=x^2 + x − 7\] Solve for x.

OpenStudy (lovelycharm):

9x+X^2

OpenStudy (lovelycharm):

9x^4

OpenStudy (john_es):

You should group the powers with the same exponent, \[x^2+10x+11-x^2-x+7=0\Rightarrow9x+18=0\Rightarrow 9x=-18\] And solve for x.

OpenStudy (lovelycharm):

i dont kind of get it to many number it kinda confused

hero (hero):

I tried showing the user an easy first step...subtract: